# Constructing Per-Shot Bitrate Ladders using Visual Information Fidelity

### v1

@@ Abstract
Video service providers need their delivery systems to be able to adapt to network conditions, user preferences, display settings, and other factors. HTTP Adaptive Streaming (HAS) provides numerous options that allow dynamic switching between different video representations to simultaneously enhance bandwidth consumption and users’ streaming experiences. Fixed bitrate ladders as employed in the past are limited in their ability to deliver high-quality visual experiences while minimizing bitrate budgets. Adaptive video streaming allows for the construction of bitrate ladders that deliver perceptually optimized visual quality to viewers under bandwidth constraints. Two common approaches to adaptation are per-title encoding and per-shot encoding. The former involves encoding each program, movie, or other content in a manner that is perceptually- and bandwidth-optimized for that content but is otherwise fixed. The latter is a more granular approach that optimizes the encoding parameters for each scene or shot (however defined) of a video content. Per-shot video encoding, as pioneered by Netflix, encodes on a per-shot basis using the Dynamic Optimizer (DO) [1]. Under the control of the VMAF perceptual video quality prediction engine, the DO delivers high-quality videos to millions of viewers at considerably reduced bitrates than per-title or fixed bitrate ladder encoding. A variety of per-title and per-shot encoding techniques have been recently proposed that seek to reduce computational overhead and to construct optimal bitrate ladders more efficiently using low-level features extracted from source videos. Here we develop a perceptually optimized method of constructing optimal per-shot bitrate and quality ladders, using an ensemble of low-level features and Visual Information Fidelity (VIF) features extracted from different scales and subbands. We compare the performance of our model, which we call VIF-ladder, against other content-adaptive bitrate ladder prediction methods, counterparts of them that we designed to construct quality ladders, a fixed bitrate ladder, and bitrate ladders constructed via exhaustive encoding using Bjontegaard delta metrics.
Index Terms:
Adaptive Streaming, Bitrate Ladder Construction, Video Processing, Gaussian Mixture Models
@@ I Introduction
A recent report on video streaming [2] stated that as of 2024, video constitutes about 74% of mobile traffic, and is predicted to increase more than 80% by 2029. While there is a noticeable increase in user-generated content (UGC) uploaded to social media platforms, the majority of video traffic can be attributed to Video-on-Demand (VoD) services offered by Netflix, Meta, YouTube, Prime Video, and others. VoD service providers deliver videos to users that are scaled and/or compressed based on their display settings, network conditions, device capabilities, available bandwidth, and buffer state. Video service providers invest substantial resources towards optimizing video compression and delivery pipelines both to decrease video transmission costs and to enhance end-user satisfaction.
@@
In recent years, HTTP Adaptive Streaming has emerged as an effective and popular standard for video content delivery, and HTTP Live Streaming (HLS) bitrate ladders [3] have become widely adopted go-to ways of adaptive bitrate streaming of all kinds of video content. HLS [3], also known as ‘one-size-fits-all’ is designed to allow adaptation to various video characteristics and network conditions. It allows the selection of different sets of video encoding parameters, based on network conditions and user preferences, to ensure a high Quality of Experience (QoE) when displayed. However, a disadvantage of HLS [3] is that the encoding settings remain the same, independent of changes in the video content, thereby not guaranteeing an optimal bitrate ladder for a given video.
@@
Recently introduced content-optimized per-shot encoding techniques [4, 5, 6] provide bitrate savings and better QoE than existing fixed bitrate ladders (HLS). Per-shot encoding techniques achieve optimal encoding by constructing a convex hull on each shot of the video. Each video title to be streamed is partitioned into shorter scenes or shots of relatively shorter durations which are encoded independently of one another. Generally, shots consist of frames that are relatively homogenous in content, hence suitable for encoding using locally fixed parameters. Each shot is encoded multiple times using a variety of bitrate-optimized encoding parameters, usually under a perceptual quality criterion, to construct a convex hull. The convex hull is where an encoding point reaches Pareto efficiency: it consists of optimal bitrate-resolutions pairs that display the highest perceptual quality at each of a set of typical bitrates. Given the extensive space of encoding settings, including spatial resolutions, bitrates, QPs or CRFs, codecs, etc, each shot is compressed under a quality criterion such as SSIM [7] multiple times when constructing the convex hull, which requires a significant amount of resources and time.
For example, assuming a single shot, and given a specific video codec, a set of $R$ resolutions, and $B$ bitrates, constructing a Pareto-front requires performing compression and quality estimation $R\times B$ times. To achieve optimal performance on transmission of a single diverse video containing multiple shots/scenes, the convex hull construction needs to be repeated many times. This, in the context of per-shot encoding, which necessitates the construction of a convex hull on each individual shot, the terms ‘shot’ and ‘video’ along with ‘convex hull’ and ‘Pareto front’ will be used interchangeably throughout.
@@
Considering the above, we present content-gnostic techniques for predicting optimal bitrate and quality ladders for various adaptive streaming services. Our proposed methods extract features from uncompressed videos at their original resolution without any compression or quality estimation, and use machine learning models to predict the quality or bitrate of a compressed video at a given resolution and bitrate or quality, respectively. With these predicted quality or bitrate values, we construct a suitable bitrate or quality ladder that yields performance comparable to a reference bitrate ladder for a given video. The reference bitrate ladder of a video is sampled from the Pareto front, which is constructed by exhaustive encoding of the video at different resolutions and rate control settings. Our proposed methods offer a distinct advantage in terms of reduced complexity, as compared to the compute-intensive process involved in constructing a reference bitrate ladder.
@@ I-A Contributions
In preliminary work [8] presented in the conference paper, we experimented with features drawn from Visual Information Fidelity (VIF) [9] to predict the quality of videos compressed by libx265, given resolution and bitrate. We utilized multiple VIF feature sets extracted from different scales and subbands of a video to predict per-shot bitrate ladders. We extracted bitrate and VMAF scores from videos, compressed using libx265 with medium preset, across eight different spatial resolutions and 33 different CRFs, to train an Extra-Trees regressor to predict the VMAF scores of compressed videos based on VIF features, bitrate, width, and height. We later used the regressor to create a per-shot bitrate ladder for a given video. We have since greatly expanded this preliminary work, making the following new contributions:
@@ I-B Paper Organization
The rest of the paper is organized in the following structure. In Section II, we discuss previous work related to the construction of per-shot bitrate ladders. Section III discusses the dataset and experimental settings we used in our simulations. We study the efficacy of combining quality-aware VIF feature sets with content-aware low-level features in Section IV. In Section V, we discuss the experimental comparisons with prior methods, as well as against their designed quality ladder predictive counterparts, and correction techniques. The paper concludes in Section VI along with future directions of research.
@@ II Related Work
As discussed in the Introduction, although a fixed bitrate ladder is designed considering various video characteristics, network conditions, resolutions, and bitrates, it is still content-independent. Despite the noteworthy performance of these ‘one-size-fits-all’ fixed bitrate ladders across a broad spectrum of videos, optimization of the encoding settings at the shot level makes it possible to further enhance the Quality of Experience (QoE) while also achieving bitrate savings. The per-shot encoding framework introduced by Netflix [4, 6, 5] performs search in encoding parameter space to determine optimal encoding settings for each shot. Over the past few years, a variety of techniques have been introduced to facilitate the construction of per-shot bitrate ladders, towards reducing reliance on exhaustive encoding methods.
@@
For example, the authors of [10], modeled PSNR as a linear function of bitrate. They extracted features including gray-level co-occurrence matrices (GLCM), temporal coherence (TC) measures, and normalized cross-correlation (NCC) to predict the coefficients of the linear function using a support vector regressor (SVR). The same authors extended their work in [11] by considering a bigger feature set and by investigating the ability of various parametric functions to fit Rate-Quality (RQ) curves. They concluded that although third-degree polynomials fit rate-quality (RQ) curves well, an exponential function was able to achieve better BD rate savings with less complexity. In [12] and [13], the authors constructed optimal Rate-PSNR curves using cross-over points. At each cross-over point, switching from a lower resolution to a higher resolution occurs given increases in bitrate. The authors modeled cross-over points between two resolutions as a pair of QPs, with a single QP defined at each resolution. Low-level features like GLCM and TC are used to predict cross-over QPs. The authors in [14] used ‘knee-points’ to construct optimal Rate-VMAF curves. The knee points were defined as QPs where a rate-quality curve has the highest curvature. They used features like the GLCM and NCC to predict the knee QPs. Instead of modeling cross-over points as QPs, the authors in [15] modeled cross-over points as bitrates. They experimented with shallow machine learning models like Extra-Trees regressor, XGBoost, and Gaussian Process regression, using features like GLCM, TC, SI, TI, and CF along with semantic-aware deep learning models like ResNet50 and VGG16, to predict cross-over bitrates between consecutive resolutions. They found the Extra-Trees regressor to perform the best.
@@
In [16], the authors modeled VMAF of a compressed video as a linear regression of bitrate and DCT-based texture energy features. They reported a good correlation against VMAF scores with the coefficients of the linear regression being allowed to vary with resolution. The authors later used this quality prediction model to construct per-shot bitrate ladders. The same authors extended their work in [17] by predicting VMAF, CRFs, and JND thresholds on compressed videos using the same DCT-based energy features, luminescence, and bitrate. The number of predicted encoding settings was further reduced during bitrate ladder construction using JND thresholds predicted using GLCM features, bitstream features, etc. The authors in [18] combined a classifier that predicts optimal resolution with a regressor that predicts cross-over bitrates, to create an ensemble aggregator that constructs bitrate ladders. The authors also used low-level spatio-temporal features like GLCM and TC. Instead of designing models to construct optimal Pareto-fronts, the authors of [19] used deep learning models to predict points on the Pareto-Front, by modeling the problem as a multi-label classification problem. The authors deployed a deep learning model with Conv-GRU units to extract spatio-temporal features on the videos before compression. Techniques like incremental learning to achieve tractable memory footprints, and transfer learning to analyze wide ranges of content complexities were used to augment training of their deep learning models.
@@
In earlier preliminary work [8], we trained an Extra-Trees regressor on multiple Visual Information Fidelity (VIF) [9] feature sets extracted over different scales and subbands of source videos. In addition to the VIF features, we used metadata like bitrates, and the widths, and heights of the compressed videos to predict the VMAF of the corresponding compressed videos. We used these models to construct per-shot bitrate ladders. One advantage of using VIF feature sets is that they are readily available when using widely deployed full-reference (FR) quality estimation model VMAF.
III Dataset and Experimental Settings
III-A Dataset
Our video data was drawn from the BVT-100 4K dataset used in [12],[13]. This dataset consists of video sequences derived from various public sources including Netflix Chimera [20], Ultra Video Group [21], Harmonic Inc. [22], SJTU [23], and AWS Elemental [24]. All of the video sequences were converted to 4:2:0 chroma subsampling, spatially cropped to UHD (3840$\times$2160 pixels), and temporally clipped to 64 frames. Each video was constrained to contain a single scene (without scene cuts). The majority of the test videos have frame rates of 60 fps and bit depths of 10 bits/pixel of luma and chroma. Fig. 1 shows sample frames of the videos comprising in the dataset.
Figure 1: Sample frames from each video taken from the BVT-100 4K dataset [13].
III-B Experiment Settings
(a) Heat-map of RQ points over 8 resolutions and 31 CRFs.
(b) Box-plot of VMAF vs spatial resolution.
(c) Box-plot of VMAF vs CRFs.
We used ffmpeg to perform compression and quality estimation. The videos were compressed using the libx265 codec with the medium preset. We used VMAF [25] to predict the perceptual qualities of the compressed videos. VMAF has been shown to exhibit higher correlations with human judgments than PSNR. We computed VMAF after upscaling the compressed video to its original resolution (3840$\times$2160). We used the Lanczos interpolation to conduct spatial upscaling and downscaling (from 2160p). We considered eight different resolutions ranging from 2160p to 288p: 3840$\times$2160, 2560$\times$1440, 1920$\times$1080, 1280$\times$720, 960$\times$540, 768$\times$432, 640$\times$360, and 512$\times$288 all of which have aspect ratio of about 16 :::: 9. We implemented constant-quality encoding by adjusting CRFs, aiming to maintain consistent visual quality while maximizing compression efficiency. We sampled libx265 CRF values ranging from 15 to 45 (inclusive).
To understand the distributions of rate-quality points for various bitrate and VMAF ranges, we plotted a heat map of rate-quality points for all videos compressed using the initial set of 8 resolutions and 31 CRFs, as shown in Fig. 2(a). To further investigate the role of each considered resolution and CRF, we plotted box plots of the distribution of VMAF for each resolution and CRF pair, as shown in Fig. 2(b) and Fig. 2(c), respectively. It may be observed that the distribution of VMAF scores was negatively skewed for higher resolutions and smaller CRFs, while VMAF scores are positively skewed for lower resolutions and larger CRFs. The majority of the rate-quality points at resolutions 640$\times$360 and 512$\times$288 and CRFs greater than 41 yielded VMAF scores less than 15. From the curves of median and mean in Fig. 2(c), it may also be observed that for large and small CRF values, the relative changes of the distribution of VMAF scores were slowly decreasing, with the relative change being the highest at intermediate CRFs.
  • Codec, Preset: libx265, medium
  • Resolutions: 3840$\times$2160, 2560$\times$1440, 1920$\times$1080, 1280$\times$720, 960$\times$540, and 768$\times$432
  • CRFs: 16 to 35 (inclusive), 35 to 41 (inclusive) with a skip of 2
  • Constraints: 15 $\leq$ VMAF $\leq$ 95
IV Method
Next, we describe the quality-aware perceptual (VIF) features and content-aware low-level features that we use to build our models.
IV-A VIF Feature Sets

### v2

@@ Abstract
Video service providers need their delivery systems to be able to adapt to network conditions, user preferences, display settings, and other factors. HTTP Adaptive Streaming (HAS) offers dynamic switching between different video representations to simultaneously enhance bandwidth consumption and users’ streaming experiences. Per-shot encoding, pioneered by Netflix, optimizes the encoding parameters on each scene or shot. The Dynamic Optimizer (DO) uses the Video Multi-Method Assessment Fusion (VMAF) perceptual video quality prediction engine to deliver high-quality videos at reduced bitrates. Here we develop a perceptually optimized method of constructing optimal per-shot bitrate and quality ladders, using an ensemble of low-level features and Visual Information Fidelity (VIF) features. During inference, our method predicts the bitrate or quality ladder of a source video without any compression or quality estimation. We compare the performance of our model against other content-adaptive bitrate ladder prediction methods, a fixed bitrate ladder, and reference bitrate ladders constructed via exhaustive encoding using Bjøntegaard-delta (BD) metrics. Our proposed method shows excellent gains in bitrate and quality against the fixed bitrate ladder and only small losses against the reference bitrate ladder, while providing significant computational advantages.
@@ I Introduction
A recent report on video streaming [Video-Streaming-Report] stated that as of 2024, video constitutes about 74% of mobile traffic, and is predicted to increase more than 80% by 2029. While there is a noticeable increase in user-generated content (UGC) uploaded to social media platforms, the majority of video traffic can be attributed to Video-on-Demand (VoD) services offered by Netflix, Meta, YouTube, Prime Video, and others. VoD service providers deliver videos to users that are scaled and/or compressed based on their display settings, network conditions, device capabilities, available bandwidth, and buffer state. Video service providers invest substantial resources towards optimizing video compression and delivery pipelines both to decrease video transmission costs and to enhance end-user satisfaction.
@@
In recent years, HTTP Adaptive Streaming has emerged as an effective and popular standard for video content delivery, and HTTP Live Streaming (HLS) bitrate ladders [Fixed-Bitrate-Ladder] have become widely adopted go-to ways of adaptive bitrate streaming of all kinds of video content. HLS [Fixed-Bitrate-Ladder], also known as ‘one-size-fits-all’ is designed to allow adaptation to various video characteristics and network conditions. It allows the selection of different sets of video encoding parameters, based on network conditions and user preferences, to ensure a high Quality of Experience (QoE) when displayed. However, a disadvantage of HLS [Fixed-Bitrate-Ladder] is that the encoding settings remain the same, independent of changes in the video content, thereby not guaranteeing an optimal bitrate ladder for a given video.
@@
Recently introduced content-optimized per-shot encoding techniques [Per-Title-Encoding, Shot-Encoding, Dynamic-Optimizer] provide bitrate savings and better QoE than existing fixed bitrate ladders (HLS). Per-shot encoding techniques achieve optimal encoding by constructing a convex hull on each shot of the video. Each video title to be streamed is partitioned into shorter scenes or shots of relatively shorter durations which are encoded independently of one another. Generally, shots consist of frames that are relatively homogenous in content, hence suitable for encoding using locally fixed parameters. Each shot is encoded multiple times using a variety of bitrate-optimized encoding parameters, usually under a perceptual quality criterion, to construct a convex hull. The convex hull is where an encoding point reaches Pareto efficiency: it consists of optimal bitrate-resolutions pairs that display the highest perceptual quality at each of a set of typical bitrates. Given the extensive space of encoding settings, including spatial resolutions, bitrates, quantization parameters (QPs), and constant rate factors (CRFs), each shot is compressed under a quality criterion such as SSIM [SSIM] multiple times when constructing the convex hull, which requires a significant amount of resources and time. For example, assuming a single shot, and given a specific video codec, a set of $R$ resolutions, and $B$ bitrates, constructing a Pareto-front requires performing compression and quality estimation $R\times B$ times. To achieve optimal performance on transmission of a single diverse video containing multiple shots/scenes, the convex hull construction needs to be repeated many times. This, in the context of per-shot encoding, which necessitates the construction of a convex hull on each individual shot, the terms ‘shot’ and ‘video’ along with ‘convex hull’ and ‘Pareto front’ will be used interchangeably throughout.
@@
Considering the above, we present content-gnostic techniques for predicting optimal bitrate and quality ladders for various adaptive streaming services. Our proposed methods extract features from uncompressed videos at their original resolution without any compression or quality estimation, and use machine learning models to predict the quality or bitrate of a compressed video at a given resolution and bitrate or quality, respectively. With these predicted quality or bitrate values, we construct a suitable bitrate or quality ladder that yields performance comparable to a reference bitrate ladder for a given video. The reference bitrate ladder of a video is sampled from the Pareto front, which is constructed by exhaustive encoding of the video at different resolutions and rate control settings. Figures 1(a) and 1(b) illustrate the training and inference flows of our methods. We will elaborate on these procedures in the subsequent sections. Our proposed methods offer a distinct advantage in terms of reduced complexity, as compared to the compute-intensive process involved in constructing a reference bitrate ladder.
(a) Proposed framework for training
(b) Proposed framework for inference
Figure 1: Proposed framework for training and inference.
@@ I-A Contributions
In preliminary work [Bitrate-Ladder-Construction-using-Visual-Information-Fidelity] presented in the conference paper, we experimented with features drawn from Visual Information Fidelity (VIF) [VIF] to predict the quality of videos compressed by libx265, given resolution and bitrate. We utilized multiple VIF feature sets extracted from different scales and subbands of a video to predict per-shot bitrate ladders. We extracted bitrate and VMAF scores from videos, compressed using libx265 with medium preset, across eight different spatial resolutions and 33 different CRFs, to train an Extra-Trees [Extra-Trees] regressor that predicts the VMAF scores of compressed videos based on VIF features, bitrate, width, and height. We later used the regressor to create a per-shot bitrate ladder for a given video. We have since greatly expanded this preliminary work, making the following new contributions:
@@ I-B Paper Organization
The rest of the paper is organized as follows. Section II discusses previous work on the construction of per-shot bitrate ladders. Section III explains the computation of quality-aware VIF feature sets and content-aware low-level features and how they are used during training and inferencing in our model. Section IV discusses the dataset and experimental settings we used in our simulations, the performances of quality and bitrate prediction regressors, their effectiveness in constructing per-shot bitrate/quality ladders, and experimental comparisons with prior methods. The paper concludes in Section V along with a discussion of future directions of research.
@@ II Related Work
As discussed in the Introduction, although a fixed bitrate ladder is designed considering various video characteristics, network conditions, resolutions, and bitrates, it is still content-independent. Despite the noteworthy performance of these ‘one-size-fits-all’ fixed bitrate ladders across a broad spectrum of videos, optimization of the encoding settings at the shot level makes it possible to further enhance the Quality of Experience (QoE) while also achieving bitrate savings. The per-shot encoding framework introduced by Netflix [Per-Title-Encoding, Dynamic-Optimizer, Shot-Encoding] performs search in encoding parameter space to determine optimal encoding settings for each shot. Over the past few years, a variety of techniques have been introduced to facilitate the construction of per-shot bitrate ladders, towards reducing reliance on exhaustive encoding methods.
@@
For example, the authors of [Predicting-Video-Rate-Distortion-Curves-using-Textural-Features], modeled PSNR as a linear function of bitrate. They extracted features including gray-level co-occurrence matrices (GLCM) [GLCM], temporal coherence (TC) [TC] measures, and normalized cross-correlation (NCC) to predict the coefficients of the linear function using a support vector regressor (SVR). The same authors extended their work in [Study-of-compression-statistics-and-prediction-of-rate-distortion-curves-for-video-texture] by considering a bigger feature set and by investigating the ability of various parametric functions to fit Rate-Quality (RQ) curves. They concluded that although third-degree polynomials fit rate-quality (RQ) curves well, an exponential function was able to achieve better Bjøntegaard-delta (BD) [BD] rate savings with less complexity. In [Content-gnostic-Bitrate-Ladder-Prediction-for-Adaptive-Video-Streaming] and [Efficient-Bitrate-Ladder-Construction-for-Content-Optimized-Adaptive-Video-Streaming], the authors constructed optimal Rate-PSNR curves using cross-over points. At each cross-over point, switching from a lower resolution to a higher resolution occurs given increases in bitrate. The authors modeled cross-over points between two resolutions as a pair of QPs, with a single QP defined at each resolution. Low-level features like GLCM and TC are used to predict cross-over QPs. The authors in [VMAF-based-Bitrate-Ladder-Estimation-for-Adaptive-Streaming] used ‘knee-points’ to construct optimal Rate-VMAF curves. The knee points were defined as QPs where a rate-quality curve has the highest curvature. They used features like the GLCM and NCC to predict the knee QPs. Instead of modeling cross-over points as QPs, the authors in [Benchmarking-Learning-based-Bitrate-Ladder-Prediction-Methods-for-Adaptive-Video-Streaming] modeled cross-over points as bitrates. They experimented with shallow machine learning models like Extra-Trees regressor [Extra-Trees], XGBoost [XGBoost], and Gaussian Processes [GP] regression, using features like GLCM, TC, spatial information (SI) [SI-TI], temporal information (TI) [SI-TI], and colorfulness (CF) [CF], as well as semantic-aware deep learning models like ResNet50 and VGG16, to predict cross-over bitrates between consecutive resolutions. They found the Extra-Trees regressor to perform the best.
@@
In [Perceptually-Aware-Per-Title-Encoding-for-Adaptive-Video-Streaming], the authors modeled VMAF of a compressed video as a linear regression of bitrate and DCT-based texture energy features. They reported a good correlation against VMAF scores with the coefficients of the linear regression being allowed to vary with resolution. The authors later used this quality prediction model to construct per-shot bitrate ladders. The same authors extended their work in [Just-Noticeable-Difference-aware-Per-Scene-Bitrate-laddering-for-Adaptive-Video-Streaming] by predicting VMAF, CRFs, and JND thresholds on compressed videos using the same DCT-based energy features, luminescence, and bitrate. The number of predicted encoding settings was further reduced during bitrate ladder construction using JND thresholds predicted using GLCM features, bitstream features, etc. The authors in [Ensemble-Learning-for-Efficient-VVC-Bitrate-Ladder-Prediction] combined a classifier that predicts optimal resolution with a regressor that predicts cross-over bitrates, to create an ensemble aggregator that constructs bitrate ladders. The authors also used low-level spatio-temporal features like GLCM and TC. Instead of designing models to construct optimal Pareto-fronts, the authors of [Convex-Hull-Prediction-for-Adaptive-Video-Streaming-by-Recurrent-Learning] used deep learning models to predict points on the Pareto-Front, by modeling the problem as a multi-label classification problem. The authors deployed a deep learning model with Conv-GRU units to extract spatio-temporal features on the videos before compression. Techniques like incremental learning to achieve tractable memory footprints, and transfer learning to analyze wide ranges of content complexities were used to augment training of their deep learning models.
@@
In earlier preliminary work [Bitrate-Ladder-Construction-using-Visual-Information-Fidelity], we trained an Extra-Trees regressor on multiple Visual Information Fidelity (VIF) [VIF] feature sets extracted over different scales and subbands of source videos. In addition to the VIF features, we used metadata like bitrates, and the widths, and heights of the compressed videos to predict the VMAF of the corresponding compressed videos. We used these models to construct per-shot bitrate ladders. One advantage of using VIF feature sets is that they are readily available when using widely deployed full-reference (FR) quality estimation model VMAF.
III Method
Next, we describe the problem formulation, training and inference procedures and the features employed in our bitrate and quality ladder model.
Figure 2: An illustration showing differences between VIF feature extraction in our method and VIF quality score estimation in full-reference video quality assessment. The block diagram is similar to a figure in [VIF]. The green and red lines in the block diagram together represent the VIF features used for quality prediction and the green lines represent the VIF features used in our method.
Figure 3: VIF Feature Extraction
III-A Problem Formulation
We formulate the problem of constructing per-shot bitrate and quality ladders as a quality and bitrate prediction problem, respectively. Unlike the fixed bitrate ladder [Fixed-Bitrate-Ladder], we aim to employ machine learning methods to construct per-shot content-gnostic bitrate and quality ladders. A convex hull or reference bitrate ladder (sampled from a convex hull) is constructed by encoding the given video across multiple resolutions and rate-control settings, then using a full-reference quality estimation model to estimate the quality of the compressed video. Here, we replace the time-consuming encoding and quality estimation procedures with machine learning models that directly predict the qualities or bitrates of videos to be compressed.
III-A1 Training
Fig. 1(a) shows the training framework of our method. We train regressors using video features extracted from a source video, along with metadata obtained by compressing and estimating the quality of the compressed video. We use VMAF [VMAF] to estimate the qualities of compressed videos. We train quality prediction regressors to predict the scaled VMAF scores ($q/100$, $q$ is VMAF) of compressed videos using video features, the logarithm of bitrate ($\log_{2}(b)$), scaled widths ($\frac{w}{3840}$), and the scaled heights ($\frac{h}{3840}$) of the compressed videos. Similarly, we train bitrate prediction regressors to predict the logarithm of bitrate ($\log_{2}(b)$) of each compressed video using video features, scaled VMAF scores ($q/100$, $q$ is VMAF), scaled widths ($\frac{w}{3840}$), and the scaled heights ($\frac{h}{3840}$) of compressed videos. It is also worth noting that, unlike regressors that predict cross-over points [Content-gnostic-Bitrate-Ladder-Prediction-for-Adaptive-Video-Streaming, Efficient-Bitrate-Ladder-Construction-for-Content-Optimized-Adaptive-Video-Streaming, Benchmarking-Learning-based-Bitrate-Ladder-Prediction-Methods-for-Adaptive-Video-Streaming], the quality or bitrate prediction regressors provide supplementary information regarding the approximate bitrates or qualities of the compressed videos. We trained multiple regressors, including Extra-Trees [Extra-Trees], XG-Boost [XGBoost], and RandomForests [Random-Forest], on each task in our experiments. Machine learning models were trained separately on each feature set/approach. We found that the Extra-Trees regressor consistently delivered the best results on all our feature sets on almost every task.
III-A2 Inference
Fig. 1(b) shows the inference framework of our proposed method. A bitrate ladder contains steps of bitrate and the corresponding optimal resolution for encoding at that bitrate step. When constructing the bitrate ladders, we selected the following bitrates (in kbps) as steps: 500, 1000, 2000, 3000, 4000, 5000, 6000, 7000, 8000, 9000, 10500, 12000, 15000. Similarly, a quality ladder contains quality steps based on the full-reference model is employed and the corresponding optimal resolution for encoding at that quality step. When constructing quality ladders, we selected the following VMAF scores as steps: 25, 35, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 92.5. To predict the per-shot bitrate ladder of a given video, we employ the quality prediction regressors to predict quality scores of videos compressed at each bitrate step across multiple resolutions. The resolution that yields the highest prediction is considered as optimal at the corresponding bitrate step. Conversely, when predicting the per-shot quality ladder, we employ bitrate prediction regressors to predict the bitrates of compressed videos at each quality step across multiple resolutions. The resolution that yields the lowest prediction is considered optimal at the corresponding quality step. In both these scenarios, we construct per-shot bitrate/quality ladders without employing compression or quality estimation modules.
III-B VIF Feature Sets